OpenStudy (anonymous):

Simplify: √96 a. 4√6 b. 16√2 c. 8√6 d. 8√2

OpenStudy (lalaly):

A

OpenStudy (dumbcow):

\[\sqrt{96} = \sqrt{16*6} = \sqrt{16}\sqrt{6} = 4\sqrt{6}\]